OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 Any Real Number No Solution 0 3

OpenStudy (anonymous):

2(x+1)=2x+5 2x+2=2x+5 *Distribute* -2x-2=-2x-2 *put like terms together. x on the left and the numbers on the right* 0=3 *it doesn't equal so there would be no solution*

OpenStudy (anonymous):

distribute again 2x+2=2x+5 Now add like terms 4x=7 Now divide x=4/7 and graph here is the graph and it'll tell you itself http://www.wolframalpha.com/input/?i=plot+x%3d4%2f7&incTime=true